The hypotenuse of a right triangle is 4 cm. longer than the one leg and 18 cm. longer than the other leg. Find the length of each side of the triangle.
Answer by fractalier(6550)

Call the shortest leg x.
The hypotenuse is x + 18.
The other leg is x + 14 since it is only 4 less than the hypotenuse.
Now apply the Pythagorean Theorem: a^2 + b^2 = c^2 and we get
x^2 + (x + 14)^2 = (x + 18)^2
x^2 + x^2 + 28x + 196 = x^2 + 36x + 324
now collecting and solving we get
x^2 - 8x - 128 = 0
(x - 16)(x + 8) = 0
x = 16 or x = -8
but distance is never negative so the short leg is 16 and the longer leg is 30. The hypotenuse is 34.